Please help me solve this equation: {{The length of a rectangle is six times its width.
If the perimeter of the rectangle is 126 yd, find its area.}}}

Answer by lwsshak3(11628)

The length of a rectangle is six times its width.
If the perimeter of the rectangle is 126 yd, find its area.
***
let x=width of rectangle
6x=length of rectangle
2*length+2*width=perimeter=126
12x+2x=126
14x=126
x=9 (width)
6x=54 (length)
Area=length*width=54*9=486 sq yd